3. Solve the following IVP: ly(0) = yo, y(0) = yo where p > 0 and k 1.2. Let k=0 Use the above power series to solve the following IVP. List the first six nonzero terms the solution the differentplease help and show steps
Answer #1

Answert Given that, initial value problem is, ③ The da2 y O) = SÓ, Ylo) - Yo eso & *#1. dez 0 de , der from © as de Loya ly-K+1 (F)Oto -K +1 -K +1 i cz (yo) -K+1 substituting in ea ② è value oakti -K+I (9)= ētk+ + (%) -Kti -K+I (9) +1 -K sy. Xyou does ty***] Fakta 9 (0)= y =) ktlink 2-Ka - 2-k -K+1 7 1K C;= Yo - (*** subs têtesting svalue in San nyo **(87* J ** +